OpenStudy (anonymous):

use scientfic notation to divide the numbers. express the answer using scientfic notation retain three decimal places (6.2*10^7)/4.15810^3)

OpenStudy (anonymous):

\[\frac{6.2*10^7}{4.158*10^3}=\frac{6.2}{4.158}\frac{10^7}{10^3}\] \[1.491*10^4\]

OpenStudy (anonymous):

When dividing: \[\frac{a^m}{a^n}=a^{m-n}\]